Suppose a sample of 200 local vendors is selected as part of a survey conducted by the Ministry of Trade and it is found that 145 of the local vendors make an average income in excess of N$ 7 600 per month. a) Suppose that a 98% confidence interval estimate of the proportion of local vendors who make an average income more than N$ 7 600 per month is required. How large a sample should be taken if the desired margin of error is 8%?

b) Suppose it has been established from past studies that not more than 65% of local vendors make an average income in excess of N$ 7 600 per month. Use the critical value approach and conduct a hypothesis test at a 5% level of significance to determine whether the proportion of local vendors who make an average income in excess of N$ 7 600 per month differs from the historical level. 


1
Expert's answer
2021-09-20T16:39:50-0400

Solution:

(a):

Let "\alpha=0.02"

Then "z_{0.02\/2}=2.33"

Margin error, ME = 0.08

Required sample size = n = "(\dfrac{z_{0.02/2}}{ME})^2\times p \times (1-p)"

"\Rightarrow n=(2.33/0.08)^2(0.73)(0.27) \ [\because p=145/200=0.73]
\\\Rightarrow n=167.193\approx167"

(b):

"H_0:p\le0.65
\\H_1:p > 0.65"

"\alpha=0.05"

Then "z_{0.05}=1.64" [Using one-tailed value]

"\hat p=145/200=0.73"

Test statistic"=\dfrac{\hat p-p}{\sqrt{\dfrac{p(1-p)}{n}}}=\dfrac{0.73-0.65}{\sqrt{\dfrac{0.65(1-0.65)}{200}}}=2.37"

Now, 2.37 > 1.64, we have sufficient evidence to reject H0 .

Thus, "p>0.65" or more than 65% of local vendors make an average income in excess of N$ 7 600 per month.